Two pucks collide on an air hockey table. Puck A has a mass of 28 g and is initially traveling in the +x direction at 7.8 m/s. Puck B has a mass of 112 g and is initially at rest. After the pucks collide, puck A moves away at an angle of 45$^\circ$ above the +x axis, while puck B travels at an angle of 42$^\circ$ below the +x axis.
Calculate puck A's final speed and puck B's final speed. Also, what fraction of kinetic energy was lost in this collision?
Answer #1
The mass of puck A is mA = 28 m = 0.028 kg
The mass of puck B is mB = 112 g = 0.112 kg
The initial speed of puck A is uA = 7.8 m/s (along positive X axis)
The initial speed of peck B is zero
The angle made by puck A with positive X axis after the collision is θA = 45o
The angle made by puck B with positive X axis after the collision is θB = -42o
Let vA and vB are the velocities of puck A and puck B respectively after the collision
Apply the law of conservation of linear momentum in Y axis
mA(0) + mB(0) = mAvAsinθA + mBvBsinθB
mAvAsinθA + mBvBsinθB = 0
(0.028 kg)vAsin(45o) + (0.112 kg)vBsin(-42o) = 0
0.0198vA - 0.0749vB = 0
vA = 3.7828vB
Apply the law of conservation of linear momentum in X axis
mAuA + mB(0) = mAvAcosθA + mBvBcosθB  
mAuA = mAvAcosθA + mBvBcosθB
(0.028 kg)(7.8 m/s) = (0.028 kg)vAcos(45o) + (0.112 kg)vBcos(-42o)  
0.2184 = 0.0198vA + 0.0832vB
0.2184 = 0.0198(3.7828vB) + 0.0832vB
vB = 1.38 m/s
Therefore
vA = 3.7828(1.38 m/s)
vA = 5.22 m/s
The initial kinetic energy of the system is Ki = 0.5mAuA2
Ki = 0.8518 J
The final kinetic energy of the system is Kf = 0.5mAvA2 + 0.5mBvB2
Kf = 0.4881 J
The loss in the kinetic energy is ΔK = 0.8518 J - 0.4881 J = 0.3637 J
The fraction of the loss in kinetic energy is
Kfraction = ΔK/Ki
Kfraction = 0.3637/0.8518
Kfraction = 0.427